# Towards Fully Autonomous Research Powered by LLMs: Case Study on Simulations

### v1

@@
Abstract
The advent of Large Language Models (LLMs) has created new opportunities for the automation of scientific research, spanning both experimental processes and computational simulations. This study explores the feasibility of constructing an autonomous simulation agent (ASA) powered by LLM, through sophisticated API integration, to automate the entire research process, from experimental design, remote upload and simulation execution, data analysis, to report compilation. Using a simulation problem of polymer chain conformations as a case study, we assessed the performance of ASAs powered by different LLMs including GPT-4-Turbo. Our findings revealed that ASA-GPT-4o achieved near-flawless execution on designated research missions, underscoring the potential of LLMs to manage complete scientific investigations autonomously. The outlined automation can be iteratively performed up to twenty cycles without human intervention, illustrating the potential of LLMs for large-scale autonomous research endeavors. Additionally, we discussed the intrinsic traits of ASAs in managing extensive tasks, focusing on self-validation mechanisms and the balance between local attention and global oversight.
@@ Introduction
In recent years, the rapid development of artificial intelligence technology, particularly the emergence of large language models (LLMs), has brought revolutionary changes to scientific research[1, 2, 3, 4, 5]. Since the launch of ChatGPT by OpenAI, LLMs such as GPT-4 and Anthropic’s Claude have demonstrated outstanding performance in natural language processing, text generation, and translation [6, 7]. These AI models not only excel in traditional text processing tasks but also show significant potential in scientific research, gradually transforming conventional research methods and providing scientists with unprecedented tools and methodologies[8, 9, 10, 11].
@@
The application of artificial intelligence (AI) in the fields of chemistry and materials engineering is continually expanding. By integrating machine learning and automated experimental technologies, faster experimental design and data analysis can be achieved, leading to the discovery of new materials and the optimization of existing material properties in a shorter timeframe [12, 13, 14]. In the domain of physical chemistry knowledge parsing, Meta’s Galactica assists researchers with tasks such as information organization, knowledge inference, and writing, showcasing the immense potential of large language models in scientific and professional domains [13]. Meanwhile, DeepMind’s AlphaFold series of algorithms enable the prediction of structures for biological molecules, from proteins and nucleic acids to larger and more complex complexes [15, 16]. In the fields of chemical synthesis and drug design, AI aids in complex chemical synthesis planning, chemical property prediction, and the generation of new materials. Integrating LLMs with specialized chemical tools and databases can extend their capabilities beyond general language tasks [17, 18, 19, 20, 21]. Additionally, some studies have highlighted the crucial role of prompt engineering in guiding models to produce accurate and useful outputs. Techniques such as Chain of Thought, ReAct, and Tree of Thoughts help structure the model’s reasoning process, reduce errors, and improve the quality of generated results [20, 22, 23, 24].
@@
Recent research shows that continuous human-AI interaction can accelerate various stages of scientific research. However, this human-led, AI-assisted model still limits the speed and depth of scientific research. Some work has explored an AI-dominated research model, such as smart laboratories (SDLs) that integrate AI, laboratory automation, and robotics to conduct autonomous experiments. These systems can automatically adjust and optimize based on intelligent analysis and decision-making support from experimental results [25, 26, 27]. The ultimate goal for future AI scientists is to achieve full automation of scientific research processes, from experimental design to data analysis and conclusion generation. This end-to-end automation is set to become a major development in chemical research and a primary objective in chemical engineering, including both experimental and simulation aspects.
@@
Random walk and self-avoiding walk models are of significant physical importance in describing the spatial configuration of polymer chains. The random walk model assumes that each ‘step’ of the chain is independent and random, while the self-avoiding walk model considers the volume exclusion effects between chain segments. Therefore, they exhibit different dynamic behaviors and scaling properties, with different scaling relationships between the mean square end-to-end distance and the number of chain segmentss N. This scaling relationship has been theoretically derived by Flory, De Gennes, and others [28, 29] and validated through simulations and experiment [30, 31]. This problem is fundamental in polymer physics and has been extensively studied. However, for AI, it still presents a novel challenge. The difficulty of predicting the scaling exponent makes it neither too easy nor prohibitively difficult, thus serving as an excellent model problem for testing AI capabilities. We prompted the AI to write simulation programs for both random walk and self-avoiding walk models, compute the mean square end-to-end distance for different chain segment numbers N, fit the results to $\left\langle R^{2}\right\rangle\propto N^{\nu}$ , determine the scaling exponent $\nu$, plot the data, and write a complete scientific report.
@@
Figure 1: A. Schematic diagram of the AI automation process for theoretical simulation research conducted in this study. The human researcher writes the research plan (RP) for the theoretical simulation and provides it to the autonomous simulation agent (ASA), which is developed by sophisticated API programming (see Method section). The ASA first generates the simulation program, then uploads it to a remote server, performs simulation calculations with different parameter conditions, collects data, and finally writes a report based on the simulation results. B. Diagram of the LLM dialogue. The human researcher only inputs the RP at the beginning, and then the ASA makes decisions based on the dialogue history, achieving a fully autonomous research mode. C. Two typical simulation problems studied in this paper: the chain conformation simulation of a random walk and the gravitational model simulation of asteroid orbit prediction (See SI).
@@
Overall, this paper presents the first fully autonomous AI-conducted simulation research process powered by LLMs. By adopting this new research paradigm, we aim to explore the potential of AI in highly complex and data-intensive scientific research, breaking traditional human limitations and advancing the automation and intelligence of scientific research.
Figure 2: Overview of RP 1-3. We designed three RPs, each containing multiple steps such as simulation, plotting, and report writing, and provided them to the ASA. RP 1 and RP 3 required the ASA to generate a complete Python program to simulate polymer chains, run the simulation locally, and produce a final report. RP 2 required the ASA to modify the provided simulation program, run the simulation on a remote server, and then generate a report. The provided simulation program (omitted in the figure) and remote server information were included in RP 2.
@@ Results
We provided RP 1 and 2 directly to the Master AI, adjusting the ASA to ensure each sub-task was handled by a distinct Subordinate AI without access to the Main AI’s or other Subordinate AIs’ conversation histories. The Main AI solely presents the AI RPs (Figure 4A) and received reports upon task completion.
We also employed GPT-4-turbo for this experiment. The results (SI-data-2) revealed that for RP 1, the Main AI effectively decomposes the mission into sub-tasks and successfully distributed them to Subordinate AIs, thereby completing RP 1. However, for RP 2, the Main AI encounters repeated failures. The primary issue lies in precise information transmission. RP 2 provides a random walk simulation program and details of a remote server, including hostname and username. Despite multiple attempts, the Main AI struggles to convey both the simulation program and server details accurately to the Subordinate AIs (Figure 4B). We attempted to instruct the Main AI to include all relevant information in the AI RPs, but this did not notably enhance the accuracy of information transmission. Additionally, we observed instances where AI RPs contained irrelevant information, potentially confusing the Subordinate AIs about the mission’s focus.
@@
RP 4 is designed for this purpose. It is nested within RP 1, instructing the Primary AI to execute the command “python AutoProg.py -s p1.txt -n i” 20 times. Each execution generates an Agent AI to fulfill RP 1 contained in p1.txt, while the Primary AI collects all generated files and conducts result analysis. The detailed content of RP 4 is depicted in Figure 4C.
@@
Figure 3: Completion Rates of Various ASAs for RP 1-3. A. Table of LLM models used in this study. B. Statistics of the number of criteria met by each ASA over 20 experiments. We established seven criteria for RP 1-3 to measure mission completion rates and counted the number of times each ASA met these criteria across 20 experiments. Some zero results are not displayed. C. ASA Scoring. Relative scores for each ASA in RP 1-3 were calculated using the EWM and TOPSIS methods. D. ASA Generated Research Report Example. Demonstrates a Word report generated by ASA-GPT-4 for RP 1, including four sections, chain conformation diagrams and a scaling relation fit plot (also see SI).
Figure 4: A. Example of Main AI’s Response. The Main AI first breaks down the overall mission into sub-tasks, then generates AI RPs beginning with ’< < < prompt’ and ending with ’end > > >’. The AI-generated RPs are then passed to subordinate AIs for execution. B. Example of Faulty and Incomplete AI RP. The black box highlights where the AI RP fails to provide necessary information such as username and password for connecting to the remote server, and the red box highlights where the AI RP includes irrelevant information. C. Overview of RP 4. The command ’python AI4SCI _VE13.py -s p1.txt -n i’ generates an Agent AI to complete the mission of RP 1, where AI4SCI _VE13.py is the API automation execution program, p1.txt is the text file containing the RP, and i denotes the $i^{th}$ execution of the command.
@@ Discussion
The experiments demonstrate that LLMs, with their robust coding ability [24, 32, 33, 34], can be directly employed in developing agents to autonomously complete complex research missions when aided by proper API programming. ASA-GPT-4o and ASA-GPT-4-turbo, notably, exhibits near-perfect completion rates for simpler missions, highlighting their potential. Multi-tier RP designs effectively handle more complex missions, though accurate information transmission remains challenging for such RPs. Enhancements in RP design and AI coordination are essential for fully harnessing AI’s potential in automating scientific research. Below, we discuss several problems existing in the current reaserch.
@@
Second, ASAs often underperform on technical aspects of tasks due to a lack of specialized domain knowledge. For example, when simulating a random walk, the algorithm needs to generate unit vectors uniformly distributed on a sphere to ensure isotropy. Although ASAs can often generate spatially random unit vectors, they fail to ensure uniform distribution on the sphere. In another example, for simulating self-avoiding walks, ASAs often incorrectly extend the random walk strategy by merely checking distances between points, resulting in high-energy state samples with large statistical errors. Correct sampling should use importance sampling to generate statistically significant conformations[35]. To avoid these issues in research, reliable solutions should be provided in the RPs, or relevant domain knowledge should be used for fine-tuning the LLMs to enhance their understanding.
@@
This phenomenon may be associated with the LLM’s propensity to depend heavily on past dialogue content. Research has documented instances where LLMs utilize context to deliver responses preferred by users, indicating an excessive reliance on previous information [36, 37, 38]. While this capability allows LLMs to adapt responses based on short dialogue histories, it can also lead to persistent errors.
@@ Conclusion
This study demonstrates the potential of LLM powered research agent ASA to autonomously conduct scientific research missions, using a polymer theory simulation problem and a more challenging asteroid-trajectory prediction problem as model problems. This process, encompassing simulation coding, remote execution, data analysis, and report generation, represents the first fully autonomous AI-directed simulation research. The notable success herein can be largely attributed to the sophisticated API programming. Agents such as ASA-GPT-4o demonstrate high completion rates for missions, a strong grasp of task requirements, and consistent alignment with overall objectives across extensive tasks. Notably, the performance exhibited in completing RP 4, entailing a hierarchy of nested tasks, surpassed expectations in a significant manner. These capabilities underscore the readiness of ASAs to independently manage long-term scientific research missions, offering a new perspective on AI applications in science as well as in chemical engineering.
Method
@@
The EWM, a well-established approach for determining the relative significance of criteria in decision-making [39], was utilized to assign weights based on data variability. This ensures criteria with higher variability, and thus more informational value, receive higher weights in the final evaluation.
TOPSIS, originally formulated by Hwang and Yoon [40], was subsequently applied. This method computes the geometric distance between each agent’s performance and both ideal and anti-ideal solutions, facilitating a comparative assessment of ASA efficacy.

### v2

@@
ABSTRACT: The advent of Large Language Models (LLMs) has created new opportunities for the automation of scientific research, spanning both experimental processes and computational simulations. This study explores the feasibility of constructing an autonomous simulation agent (ASA) powered by LLM, through sophisticated API integration, to automate the entire research process, from experimental design, remote upload and simulation execution, data analysis, to report compilation. Using a simulation problem of polymer chain conformations as a case study, we assessed the performance of ASAs powered by different LLMs including GPT-4-Turbo. Our findings revealed that ASA-GPT-4o achieved near-flawless execution on designated research missions, underscoring the potential of LLMs to manage complete scientific investigations autonomously. The outlined automation can be iteratively performed up to twenty cycles without human intervention, illustrating the potential of LLMs for large-scale autonomous research endeavors. Additionally, we discussed the intrinsic traits of ASAs in managing extensive tasks, focusing on self-validation mechanisms and the balance between local attention and global oversight. KEYWORDS: autonomous research, simulation, LLM, polymer physics
footnotetext: a The State Key Laboratory of Molecular Engineering of Polymers, The Research Center of AI for Polymer Science Department of Macromolecular Science, Fudan University, Shanghai 200433, China; E-mail: $\ast$ lijf@fudan.edu.cn
@@ INTRODUCTION
In recent years, the rapid development of artificial intelligence technology, particularly the emergence of large language models (LLMs), has brought revolutionary changes to scientific research1, 2, 3, 4, 5. Since the launch of ChatGPT by OpenAI, LLMs such as GPT-4 and Anthropic’s Claude have demonstrated outstanding performance in natural language processing, text generation, and translation 6, 7. These AI models not only excel in traditional text processing tasks but also show significant potential in scientific research, gradually transforming conventional research methods and providing scientists with unprecedented tools and methodologies8, 9, 10, 11.
@@
The application of artificial intelligence (AI) in the fields of chemistry and materials engineering is continually expanding. By integrating machine learning and automated experimental technologies, faster experimental design and data analysis can be achieved, leading to the discovery of new materials and the optimization of existing material properties in a shorter timeframe 12, 13, 14. In the domain of physical chemistry knowledge parsing, Meta’s Galactica assists researchers with tasks such as information organization, knowledge inference, and writing, showcasing the immense potential of large language models in scientific and professional domains 13. Meanwhile, DeepMind’s AlphaFold series of algorithms enable the prediction of structures for biological molecules, from proteins and nucleic acids to larger and more complex complexes 15, 16. In the fields of chemical synthesis and drug design, AI aids in complex chemical synthesis planning, chemical property prediction, and the generation of new materials. Integrating LLMs with specialized chemical tools and databases can extend their capabilities beyond general language tasks 17, 18, 19, 20, 21. Additionally, some studies have highlighted the crucial role of prompt engineering in guiding models to produce accurate and useful outputs. Techniques such as Chain of Thought, ReAct, and Tree of Thoughts help structure the model’s reasoning process, reduce errors, and improve the quality of generated results 20, 22, 23, 24.
@@
Recent research shows that continuous human-AI interaction can accelerate various stages of scientific research. However, this human-led, AI-assisted model still limits the speed and depth of scientific research. Some work has explored an AI-dominated research model, such as smart laboratories (SDLs) that integrate AI, laboratory automation, and robotics to conduct autonomous experiments. These systems can automatically adjust and optimize based on intelligent analysis and decision-making support from experimental results 25, 26, 27. The ultimate goal for future AI scientists is to achieve full automation of scientific research processes, from experimental design to data analysis and conclusion generation. This end-to-end automation is set to become a major development in chemical research and a primary objective in chemical engineering, including both experimental and simulation aspects.
@@
Random walk and self-avoiding walk models are of significant physical importance in describing the spatial configuration of polymer chains. The random walk model assumes that each ‘step’ of the chain is independent and random, while the self-avoiding walk model considers the volume exclusion effects between chain segments. Therefore, they exhibit different dynamic behaviors and scaling properties, with different scaling relationships between the mean square end-to-end distance and the number of chain segmentss N. This scaling relationship has been theoretically derived by Flory, De Gennes, and others 28, 29 and validated through simulations and experiment 30, 31. This problem is fundamental in polymer physics and has been extensively studied. However, for AI, it still presents a novel challenge. The difficulty of predicting the scaling exponent makes it neither too easy nor prohibitively difficult, thus serving as an excellent model problem for testing AI capabilities. We prompted the AI to write simulation programs for both random walk and self-avoiding walk models, compute the mean square end-to-end distance for different chain segment numbers N, fit the results to $\left\langle R^{2}\right\rangle\propto N^{\nu}$ , determine the scaling exponent $\nu$, plot the data, and write a complete scientific report.
@@
Fig.  1: A. Schematic diagram of the AI automation process for theoretical simulation research conducted in this study. The human researcher writes the research plan (RP) for the theoretical simulation and provides it to the autonomous simulation agent (ASA), which is developed by sophisticated API programming (see Method section). The ASA first generates the simulation program, then uploads it to a remote server, performs simulation calculations with different parameter conditions, collects data, and finally writes a report based on the simulation results. B. Diagram of the ASA dialogue. The human researcher only inputs the RP at the beginning, and then the ASA makes decisions based on the dialogue history, achieving a fully autonomous research mode. C. Two typical simulation problems studied in this paper: the chain conformation simulation of a random walk and the gravitational model simulation of asteroid orbit prediction (See SI).
@@
Overall, this paper presents the first fully autonomous AI-conducted simulation research process powered by LLMs. As demonstrated in the demo video, simply submitting a research plan file via the command ’do AI4S_prj1.txt’ to ASA allows it to execute all subsequent tasks without any human intervention; ASA can even automatically upload programs to a remote server and perform a series of simulations—-something that cannot be achieved solely through interactions with a web-based LLM. By adopting this new research paradigm, we aim to explore the potential of AI in highly complex and data-intensive scientific research, breaking traditional human limitations and advancing the automation and intelligence of scientific research.
@@ Results
Fig.  2: Overview of RP 1-3. We designed three RPs related to polymer chain simulation, each containing multiple steps such as simulation, plotting, and report writing, and provided them to the ASA. RP 1 and RP 3 required the ASA to generate a complete Python program to simulate polymer chains, run the simulation locally, and produce a final report. RP 2 required the ASA to modify the provided simulation program, run the simulation on a remote server, and then generate a report. The provided simulation program (omitted in the figure) and remote server information were included in RP 2.
Fig.  3: Completion Rates of Various ASAs for RP 1-3. A. Table of LLM models used in this study. B. Statistics of the number of criteria met by each ASA over 20 experiments. We established seven criteria for RP 1-3 to measure mission completion rates and counted the number of times each ASA met these criteria across 20 experiments. Some zero results are not displayed. C. ASA Scoring. Relative scores for each ASA in RP 1-3 were calculated using the EWM and TOPSIS methods. D. ASA Generated Research Report Example. Demonstrates a Word report generated by ASA-GPT-4 for RP 1, including four sections, chain conformation diagrams and a scaling relation fit plot (also see SI).
@@
We provided RP 1 and 2 directly to the Master AI, adjusting the ASA to ensure each sub-task was handled by a distinct Subordinate AI without access to the Main AI’s or other Subordinate AIs’ conversation histories. The Main AI solely presents the AI RPs (Figure S1A) and received reports upon task completion.
We also employed GPT-4-turbo for this experiment. The results (SI-data-2) revealed that for RP 1, the Main AI effectively decomposes the mission into sub-tasks and successfully distributed them to Subordinate AIs, thereby completing RP 1. However, for RP 2, the Main AI encounters repeated failures. The primary issue lies in precise information transmission. RP 2 provides a random walk simulation program and details of a remote server, including hostname and username. Despite multiple attempts, the Main AI struggles to convey both the simulation program and server details accurately to the Subordinate AIs (Figure S1B). We attempted to instruct the Main AI to include all relevant information in the AI RPs, but this did not notably enhance the accuracy of information transmission. Additionally, we observed instances where AI RPs contained irrelevant information, potentially confusing the Subordinate AIs about the mission’s focus.
@@
RP 4 is designed for this purpose. It is nested within RP 1, instructing the Primary AI to execute the command “python AutoProg.py -s p1.txt -n i” 20 times. Each execution generates an Agent AI to fulfill RP 1 contained in p1.txt, while the Primary AI collects all generated files and conducts result analysis. The detailed content of RP 4 is depicted in Figure S1C.
@@ Discussion
The experiments demonstrate that LLMs, with their robust coding ability 24, 32, 33, 34, can be directly employed in developing agents to autonomously complete complex research missions when aided by proper API programming. ASA-GPT-4o and ASA-GPT-4-turbo, notably, exhibits near-perfect completion rates for simpler missions, highlighting their potential. Multi-tier RP designs effectively handle more complex missions, though accurate information transmission remains challenging for such RPs. We emphasize that these simulation tasks, particularly those involving remote server execution, cannot be accomplished through interactions with web-based LLMs alone.
Enhancements in RP design and AI coordination are essential for fully harnessing AI’s potential in automating scientific research. Below, we discuss the insights and existing problems identified in the current research.
Recommended Practices for Fully Autonomous Research
ASA achieves fully autonomous research through continuous interaction with the LLM according to a specific logical design guided by the research plan (RP) provided by the researchers. We have identified several key practices in ASA logic design and RP design to enhance execution success rates.
  • Structure agent responses for efficient code and task extraction. Require the LLM to produce structured outputs, enabling ASA to accurately extract Python code, subtasks, and progress indicators. This improves code execution success rates and ensures smooth mission progression.
  • Incorporate agent-led code review and debugging. While the LLM is powerful in code generation, it may struggle with complex programs on the first attempt. By integrating self-review and debugging process, the LLM can generate correct and task-compliant code within several dialogue rounds.
  • Ensure key outputs at each subtask step for continuity. Long missions often need to be divided into subtasks. To maintain coherence, prompt the LLM to consider previous outputs, generate key results for each step and input key results into the proceeding round of dialogue, facilitating the continuation of subsequent tasks and allowing the agent to monitor overall progress.
  • Provide detailed information for complex operations like remote server connections. Providing sufficient information in RP can improve mission success rates. In addition to necessary numeric/textual information, sometimes offering specific methods can guide the LLM to write correct programs more efficiently.
In summary, while the LLM has strong generative capabilities, its unpredictable outputs can complicate automation. ASA, based on Python’s AutoProg, though less intelligent, is reliable when logic is well-designed. By effectively structuring logic, ASA can oversee the LLM’s output, enhancing process reliability and achieving fully autonomous research.
Fig.  4: Recommended practices for fully autonomous research. By conducting iterative dialogues with the LLM adhering to a specific design logic, the success rate and efficiency of the automated process are enhanced. The diagrams on the left and right illustrate scenarios without and with the implementation of these strategies, respectively.
@@
Second, ASAs often underperform on technical aspects of tasks due to a lack of specialized domain knowledge. For example, when simulating a random walk, the algorithm needs to generate unit vectors uniformly distributed on a sphere to ensure isotropy. Although ASAs can often generate spatially random unit vectors, they fail to ensure uniform distribution on the sphere. In another example, for simulating self-avoiding walks, ASAs often incorrectly extend the random walk strategy by merely checking distances between points, resulting in high-energy state samples with large statistical errors. Correct sampling should use importance sampling to generate statistically significant conformations35. To avoid these issues in research, reliable solutions should be provided in the RPs, or relevant domain knowledge should be used for fine-tuning the LLMs to enhance their understanding.
@@
This phenomenon may be associated with the LLM’s propensity to depend heavily on past dialogue content. Research has documented instances where LLMs utilize context to deliver responses preferred by users, indicating an excessive reliance on previous information 36, 37, 38. While this capability allows LLMs to adapt responses based on short dialogue histories, it can also lead to persistent errors.
@@ Conclusion
This study demonstrates the potential of LLM powered research agent ASA to autonomously conduct scientific research missions, using a polymer theory simulation problem and a more challenging asteroid-trajectory prediction problem as model problems. This process, encompassing simulation coding, remote execution, data analysis, and report generation, represents the first fully autonomous AI-directed simulation research: once the research plan is made, only a single command (e.g., ‘do prj.txt’) is required to initiate and complete the process. The notable success herein can be largely attributed to the sophisticated API programming. Agents such as ASA-GPT-4o demonstrate high completion rates for missions, a strong grasp of task requirements, and consistent alignment with overall objectives across extensive tasks. Notably, the performance exhibited in completing RP 4, entailing a hierarchy of nested tasks, surpassed expectations in a significant manner. These capabilities underscore the readiness of ASAs to independently manage long-term scientific research missions, offering a new perspective on AI applications in science as well as in chemical engineering.
Methods
@@
The EWM, a well-established approach for determining the relative significance of criteria in decision-making 39, was utilized to assign weights based on data variability. This ensures criteria with higher variability, and thus more informational value, receive higher weights in the final evaluation.
TOPSIS, originally formulated by Hwang and Yoon 40, was subsequently applied. This method computes the geometric distance between each agent’s performance and both ideal and anti-ideal solutions, facilitating a comparative assessment of ASA efficacy.
@@
ASSOCIATED CONTENT
Supporting Information
The following Supporting Information is available free of charge.
Partial results for gravitational simulation mission along with examples of reports written by ASAs (SI.pdf); partial AutoProgs and experimental result files for RP 1-3, main AI generating RP, and Multi-tier human RP (Some information has been obscured; readers can replace it with their own information and follow the instructions to run the AutoProgs) (SI-data-1.zip and SI-data-2.zip); a demo video demonstrating the operation process of the automatic research system (SI-video.mp4).
AUTHOR INFORMATION
Corresponding Author
Jianfeng Li - The State Key Laboratory of Molecular Engineering of Polymers, The Research Center of AI for Polymer Science
Department of Macromolecular Science, Fudan University, Shanghai 200433, China; Email: lijf@fudan.edu.cn
Authors
Zhihan Liu - The State Key Laboratory of Molecular Engineering of Polymers, The Research Center of AI for Polymer Science
Department of Macromolecular Science, Fudan University, Shanghai 200433, China
Yubo Chai - The State Key Laboratory of Molecular Engineering of Polymers, The Research Center of AI for Polymer Science
Department of Macromolecular Science, Fudan University, Shanghai 200433, China
Author Contributions
Z.H.L. was the primary author of the paper. Y.B.C. was responsible for part of the data analysis and statistics. J.F.L. originally initiated and supervised the project.
Notes
The authors declare no competing financial interest.